OpenStudy (wesdg1978):

For the curve f(x)=2x^2-2x+1, how do you find an equation of the tangent line that has slope 2?

OpenStudy (anonymous):

get the derivative make the derivative equal to 2 and solve for x

OpenStudy (wesdg1978):

You mean, take 4x-2 and make it 4x=2 and x=2?

OpenStudy (wesdg1978):

Or 4x-2=2?

OpenStudy (anonymous):

\[\frac{d}{dx}2x^2 - 2x + 1\]\[\frac{d}{dx}=4x -2\]\[4(2) -2 = ?\]

OpenStudy (wesdg1978):

6

OpenStudy (wesdg1978):

But now what?

OpenStudy (anonymous):

Use point slope form to find y1

OpenStudy (wesdg1978):

I don't have any points to plug into point slope form or am I missing something?

OpenStudy (anonymous):

you're missing something

OpenStudy (wesdg1978):

I have slope 2 and whatever 6 is after I put 2 in for x . Now I'm lost

OpenStudy (anonymous):

The 6 is the X coordinate that we're dealing with

OpenStudy (wesdg1978):

Well, I can get to y-ysub1=2x-12 with what I have but then how do I find ysub1?

OpenStudy (anonymous):

Use an initial condition to find the Y intercept

OpenStudy (wesdg1978):

I don't understand what that means, I've never heard of an initial condition.

OpenStudy (zzr0ck3r):

say we have f(x) = x+2+c and we have initial condition that f(3) =2 we can use this to solve for c 2=3+2+c c=-3

OpenStudy (zzr0ck3r):

an initial condition is just a given point

OpenStudy (wesdg1978):

OK, that makes sense, the only given point was slope 2 though

OpenStudy (wesdg1978):

So I still don't understand how to find it.

OpenStudy (zzr0ck3r):

are you sure you were not given more? tangent lines are at a point.

OpenStudy (wesdg1978):

This is the question here: For the curve f(x)=2x^2-2x+1, find an equation of the tangent line that has slope 2?

OpenStudy (wesdg1978):

That's the only information given.

OpenStudy (zzr0ck3r):

4x-2=2 so x = 1 only one spot on the curve has a derivative of 2 and thus only one line so we know the derivative at x=1 is 2 so plug 1 in for x in the original equation and you get 1 so your point is (1,1)

OpenStudy (zzr0ck3r):

y=2x+b with point (1,1) 1=2+b so b=-1 so your equation is y=2x-1 this is the tangent line, to your curve at the point (1,1)

OpenStudy (zzr0ck3r):

and is the only tangent line to your curve with slope 2

OpenStudy (zzr0ck3r):

understand?

OpenStudy (wesdg1978):

I'm going through it in my head step by step. So, I don't need the x=6 or the ysub1 to find this equation then?

OpenStudy (zzr0ck3r):

I don't know why @hobbs987 wrote 4(2)-2=? f'(x) has slope as its range, the domain is still "position"

OpenStudy (zzr0ck3r):

so f'(2) makes no since in this context.

OpenStudy (wesdg1978):

OK, I think I've got it now. Thanks for your help!

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

the key to this, is your function only has 2 as its derivative at one spot, namely (1,1). if it had more than one place where its derivative was 2, then we would have more than one possible tangent line, and thus the answer would just be y=2x+b. But with only one spot (1,1) we can say for sure that b=-3

OpenStudy (wesdg1978):

Wait, why is b=-3?

OpenStudy (zzr0ck3r):

typo b=-1

OpenStudy (wesdg1978):

OK and when you say that it only has 2 as its derivative, your talking about slope 2 right?